An ambulance left the hospital and drove south. 3 hours later a police car left driving 42 k/h faster attempting to catch up to the ambulance. The officer caught up after 2 hours. Find the average speed of the ambulance.
Answer by ankor@dixie-net.com(22740)

An ambulance left the hospital and drove south.
3 hours later a police car left driving 42 k/h faster attempting to catch up to the ambulance.
The officer caught up after 2 hours. Find the average speed of the ambulance.
:
Let a = the speed of the ambulance
then
(a+42) = the speed of the police car
:
From the information given, we know the travel time of the ambulance = 5 hrs
and the Cop = 2 hrs
:
When the cop catches the amb, they will have traveled the same distance
Write a distance equation
dist = speed * time
:
5a = 2(a+42)
5a = 2a + 84
5a - 2a = 84
3a = 84
a = 84/3
a = 28 mph speed of the ambulance
:
:
Confirm this by finding the distances, should be equal
5(28) = 140 km
2(28+42) = 140 km